# The Optimality of Blocking Designs in Equally and Unequally Allocated Randomized Experiments with General Response

### v3

@@ A.1 Estimator Computation
$=$ $\frac{1}{n_{T}}\frac{(\boldsymbol{1}_{n}+\boldsymbol{W})^{\top}}{2}\boldsymbol{Y}-\frac{1}{n_{C}}\frac{(\boldsymbol{1}_{n}-\boldsymbol{W})^{\top}}{2}\boldsymbol{Y}$
$=$ $\frac{1}{2}\left(\frac{(\boldsymbol{1}_{n}+\boldsymbol{W})}{n_{T}}-\frac{(\boldsymbol{1}_{n}-\boldsymbol{W})}{n_{C}}\right)^{\top}\boldsymbol{Y}$
We now substitute for $\boldsymbol{Y}$ from Equation 1. To simplify the notation, let $\boldsymbol{v}:=\boldsymbol{y}_{T}+\boldsymbol{y}_{C}$ which is slightly different than the notation in the main text, $v_{0}:=\boldsymbol{1}_{2n}\boldsymbol{v}$, $y_{T,0}:=\boldsymbol{1}_{n}^{\top}\boldsymbol{y}_{T}$ and $y_{C,0}:=\boldsymbol{1}_{2n}^{\top}\boldsymbol{y}_{C}$.
$\hat{\tau}$ $=$ $\frac{1}{4}\left(\frac{(\boldsymbol{1}_{n}^{\top}+\boldsymbol{W})^{\top}}{n_{T}}-\frac{(\boldsymbol{1}_{n}^{\top}-\boldsymbol{W}^{\top})}{n_{C}}\right)\left(\boldsymbol{v}+\text{Diag}\left[\boldsymbol{W}\right](\boldsymbol{y}_{T}-\boldsymbol{y}_{C})\right)$
$=$ $\frac{1}{4}\left(\frac{(v_{0}+\boldsymbol{W}^{\top}\boldsymbol{v})}{n_{T}}-\frac{(v_{0}-\boldsymbol{W}^{\top}\boldsymbol{v})}{n_{C}}+\left(\frac{\boldsymbol{1}_{n}^{\top}+\boldsymbol{W}^{\top}}{n_{T}}-\frac{\boldsymbol{1}_{n}^{\top}-\boldsymbol{W}^{\top}}{n_{C}}\right)\text{Diag}\left[\boldsymbol{W}\right](\boldsymbol{y}_{T}-\boldsymbol{y}_{C})\right)$
$=$ $\frac{1}{4}\left(\frac{(v_{0}+\boldsymbol{W}^{\top}\boldsymbol{v})}{n_{T}}-\frac{(v_{0}-\boldsymbol{W}^{\top}\boldsymbol{v})}{n_{C}}+\left(\frac{\boldsymbol{W}^{\top}+\boldsymbol{1}_{n}^{\top}}{n_{T}}-\frac{\boldsymbol{W}^{\top}-\boldsymbol{1}_{n}^{\top}}{n_{C}}\right)(\boldsymbol{y}_{T}-\boldsymbol{y}_{C})\right)$
@@
$=$ $\frac{1}{2}\left(\boldsymbol{1}_{2n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}-\frac{\boldsymbol{y}_{C}}{n_{C}}\right)+\boldsymbol{W}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}+\frac{\boldsymbol{y}_{C}}{n_{C}}\right)\right)=\frac{1}{2n}\left(\boldsymbol{1}_{2n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{r}-\frac{\boldsymbol{y}_{C}}{\tilde{r}}\right)+\boldsymbol{W}^{\top}\left(\frac{\boldsymbol{y}_{T}}{r}+\frac{\boldsymbol{y}_{C}}{\tilde{r}}\right)\right)$
@@ A.2 Unbiasedness of the Estimator
$\mathbb{E}_{\,\boldsymbol{W}}\left[\hat{\tau}\right]$ $=$ $\frac{1}{2}\left(\boldsymbol{1}_{n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}-\frac{\boldsymbol{y}_{C}}{n_{C}}\right)+\frac{n_{T}-n_{C}}{2n}\boldsymbol{1}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}+\frac{\boldsymbol{y}_{C}}{n_{C}}\right)\right)$
$=$ $\frac{1}{2}\boldsymbol{1}_{n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}-\frac{\boldsymbol{y}_{C}}{n_{C}}+\frac{n_{T}-n_{C}}{2n}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}+\frac{\boldsymbol{y}_{C}}{n_{C}}\right)\right)$
$=$ $\frac{1}{2}\boldsymbol{1}_{n}^{\top}\left(\boldsymbol{y}_{T}\left(\frac{1}{n_{T}}+\frac{n_{T}-n_{C}}{2nn_{T}}\right)+\boldsymbol{y}_{C}\left(-\frac{1}{n_{C}}+\frac{n_{T}-n_{C}}{2nn_{C}}\right)\right)$
$=$ $\frac{1}{2}\boldsymbol{1}_{n}^{\top}\left(\boldsymbol{y}_{T}\frac{n+n_{T}-n_{C}}{2nn_{T}}+\boldsymbol{y}_{C}\frac{-n+n_{T}-n_{C}}{2nn_{C}}\right)$
$=$ $\frac{1}{2}\boldsymbol{1}_{n}^{\top}\left(\boldsymbol{y}_{T}\frac{2n_{T}}{2nn_{T}}+\boldsymbol{y}_{C}\frac{-2n_{C}}{2nn_{C}}\right)$
$=$ $\frac{1}{2n}\boldsymbol{1}_{n}^{\top}\left(\boldsymbol{y}_{T}-\boldsymbol{y}_{C}\right)=\tau~{}~{}\blacksquare$
@@
$\mathbb{V}\text{ar}\left[W_{i}\right]$ $=$ $1-b=\frac{n_{T}n_{C}}{n^{2}}=r\tilde{r}$ (25)
@@ A.7 Variance of the MSE
$=\sum_{i=1}^{n}\sigma_{i\,,i}^{2}\mathbb{E}\left[Z_{i}^{4}\right]+\sum_{i=j\neq k=\ell}\sigma_{i\,,i}\sigma_{k\,,k}\rho_{i}\rho_{k}+\sum_{i=k\neq j=\ell}\sigma_{i\,,j}\sigma_{i\,,j}\rho_{i}\rho_{j}+\sum_{i=\ell\neq j=k}\sigma_{i\,,j}\sigma_{j\,,i}\rho_{i}\rho_{j}.$
@@
$=$ $\boldsymbol{r}^{2}\tilde{r}^{2}\sum_{i=1}^{n}\mathbb{E}\left[Z_{i}^{4}\right]+r^{2}\tilde{r}^{2}\sum_{i=j\neq k=\ell}\rho_{i}\rho_{k}+\sum_{i=k\neq j=\ell}\sigma_{i\,,j}^{2}\rho_{i}\rho_{j}+\sum_{i=\ell\neq j=k}\sigma_{i\,,j}^{2}\rho_{i}\rho_{j}$ (35)
$=$ $r^{2}\tilde{r}^{2}\left(\sum_{i=1}^{n}\mathbb{E}\left[Z_{i}^{4}\right]+\sum_{i\neq j}\rho_{i}\rho_{j}\right)+2\sum_{i\neq j}\sigma_{i\,,j}^{2}\rho_{i}\rho_{j}.$
@@
$[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}]_{i,i}=\left[\sigma_{i\,,1}\rho_{1}\ldots\sigma_{i\,,2n}\rho_{2n}\right]\left[\sigma_{1\,,j}\rho_{j}\ldots\sigma_{2n\,,j}\rho_{j}\right]^{\top}=\sum_{j=1}^{n}(\sigma_{i\,,j}\rho_{j})(\sigma_{j\,,i}\rho_{i})=\sum_{j=1}^{n}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}$
@@
$\footnotesize 2\,\text{tr}\left[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}\right]=2\sum_{i=1}^{n}\sum_{j=1}^{n}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}=2\sum_{i\neq j}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}+2\sum_{i=j}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}=2\sum_{i\neq j}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}+2r^{2}\tilde{r}^{2}\sum_{i=1}^{n}\rho_{i}^{2}$ (36)
@@
$\mathbb{E}\left[(\boldsymbol{Z}^{\top}\boldsymbol{\Sigma}_{W}\boldsymbol{Z})^{2}\right]=r^{2}\tilde{r}^{2}\left(\sum_{i=1}^{n}\mathbb{E}\left[Z_{i}^{4}\right]+\sum_{i\neq j}\rho_{i}\rho_{j}-2\sum_{i=1}^{n}\rho_{i}^{2}\right)+2\,\text{tr}\left[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}\right].$
@@
$\mathbb{V}\text{ar}_{\boldsymbol{Z}}\left[\boldsymbol{Z}^{\top}\boldsymbol{\Sigma}_{W}\boldsymbol{Z}\right]=r^{2}\tilde{r}^{2}\left(\sum_{i=1}^{n}\mathbb{E}\left[Z_{i}^{4}\right]+\sum_{i\neq j}\rho_{i}\rho_{j}-2\sum_{i=1}^{n}\rho_{i}^{2}-\text{tr}\left[\boldsymbol{\Sigma}_{Z}\right]^{2}\right)+2\,\text{tr}\left[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}\right].$
Substituting the expansion of $\text{tr}\left[\boldsymbol{\Sigma}_{Z}\right]^{2}=\sum_{i\neq j}\rho_{i}\rho_{j}+\sum_{i=1}^{n}\rho_{i}^{2}$ we arrive at
$\mathbb{V}\text{ar}_{\boldsymbol{Z}}\left[\boldsymbol{Z}^{\top}\boldsymbol{\Sigma}_{W}\boldsymbol{Z}\right]$ $=$ $r^{2}\tilde{r}^{2}\left(\sum_{i=1}^{n}\mathbb{E}\left[Z_{i}^{4}\right]-3\sum_{i=1}^{n}\rho_{i}^{2}\right)+2\,\text{tr}\left[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}\right]$ (37)

### v4

@@ A.1 Estimator Computation
$=$ $\frac{1}{n_{T}}\frac{(\boldsymbol{1}_{2n}+\boldsymbol{W})^{\top}}{2}\boldsymbol{Y}-\frac{1}{n_{C}}\frac{(\boldsymbol{1}_{2n}-\boldsymbol{W})^{\top}}{2}\boldsymbol{Y}$
$=$ $\frac{1}{2}\left(\frac{(\boldsymbol{1}_{2n}+\boldsymbol{W})}{n_{T}}-\frac{(\boldsymbol{1}_{2n}-\boldsymbol{W})}{n_{C}}\right)^{\top}\boldsymbol{Y}$
We now substitute for $\boldsymbol{Y}$ from Equation 1. To simplify the notation, let $\boldsymbol{v}:=\boldsymbol{y}_{T}+\boldsymbol{y}_{C}$ which is slightly different than the notation in the main text, $v_{0}:=\boldsymbol{1}_{2n}\boldsymbol{v}$, $y_{T,0}:=\boldsymbol{1}_{2n}^{\top}\boldsymbol{y}_{T}$ and $y_{C,0}:=\boldsymbol{1}_{2n}^{\top}\boldsymbol{y}_{C}$.
$\hat{\tau}$ $=$ $\frac{1}{4}\left(\frac{(\boldsymbol{1}_{2n}^{\top}+\boldsymbol{W})^{\top}}{n_{T}}-\frac{(\boldsymbol{1}_{2n}^{\top}-\boldsymbol{W}^{\top})}{n_{C}}\right)\left(\boldsymbol{v}+\text{Diag}\left[\boldsymbol{W}\right](\boldsymbol{y}_{T}-\boldsymbol{y}_{C})\right)$
$=$ $\frac{1}{4}\left(\frac{(v_{0}+\boldsymbol{W}^{\top}\boldsymbol{v})}{n_{T}}-\frac{(v_{0}-\boldsymbol{W}^{\top}\boldsymbol{v})}{n_{C}}+\left(\frac{\boldsymbol{1}_{2n}^{\top}+\boldsymbol{W}^{\top}}{n_{T}}-\frac{\boldsymbol{1}_{2n}^{\top}-\boldsymbol{W}^{\top}}{n_{C}}\right)\text{Diag}\left[\boldsymbol{W}\right](\boldsymbol{y}_{T}-\boldsymbol{y}_{C})\right)$
$=$ $\frac{1}{4}\left(\frac{(v_{0}+\boldsymbol{W}^{\top}\boldsymbol{v})}{n_{T}}-\frac{(v_{0}-\boldsymbol{W}^{\top}\boldsymbol{v})}{n_{C}}+\left(\frac{\boldsymbol{W}^{\top}+\boldsymbol{1}_{2n}^{\top}}{n_{T}}-\frac{\boldsymbol{W}^{\top}-\boldsymbol{1}_{2n}^{\top}}{n_{C}}\right)(\boldsymbol{y}_{T}-\boldsymbol{y}_{C})\right)$
@@
$=$ $\frac{1}{2}\left(\boldsymbol{1}_{2n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}-\frac{\boldsymbol{y}_{C}}{n_{C}}\right)+\boldsymbol{W}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}+\frac{\boldsymbol{y}_{C}}{n_{C}}\right)\right)$
$=$ $\frac{1}{2n}\left(\boldsymbol{1}_{2n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{r}-\frac{\boldsymbol{y}_{C}}{\tilde{r}}\right)+\boldsymbol{W}^{\top}\left(\frac{\boldsymbol{y}_{T}}{r}+\frac{\boldsymbol{y}_{C}}{\tilde{r}}\right)\right)$
@@ A.2 Unbiasedness of the Estimator
$\mathbb{E}_{\,\boldsymbol{W}}\left[\hat{\tau}\right]$ $=$ $\frac{1}{2}\left(\boldsymbol{1}_{2n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}-\frac{\boldsymbol{y}_{C}}{n_{C}}\right)+\frac{n_{T}-n_{C}}{2n}\boldsymbol{1}_{2n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}+\frac{\boldsymbol{y}_{C}}{n_{C}}\right)\right)$
$=$ $\frac{1}{2}\boldsymbol{1}_{2n}^{\top}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}-\frac{\boldsymbol{y}_{C}}{n_{C}}+\frac{n_{T}-n_{C}}{2n}\left(\frac{\boldsymbol{y}_{T}}{n_{T}}+\frac{\boldsymbol{y}_{C}}{n_{C}}\right)\right)$
$=$ $\frac{1}{2}\boldsymbol{1}_{2n}^{\top}\left(\boldsymbol{y}_{T}\left(\frac{1}{n_{T}}+\frac{n_{T}-n_{C}}{2nn_{T}}\right)+\boldsymbol{y}_{C}\left(-\frac{1}{n_{C}}+\frac{n_{T}-n_{C}}{2nn_{C}}\right)\right)$
$=$ $\frac{1}{2}\boldsymbol{1}_{2n}^{\top}\left(\boldsymbol{y}_{T}\frac{n+n_{T}-n_{C}}{2nn_{T}}+\boldsymbol{y}_{C}\frac{-n+n_{T}-n_{C}}{2nn_{C}}\right)$
$=$ $\frac{1}{2}\boldsymbol{1}_{2n}^{\top}\left(\boldsymbol{y}_{T}\frac{2n_{T}}{2nn_{T}}+\boldsymbol{y}_{C}\frac{-2n_{C}}{2nn_{C}}\right)$
$=$ $\frac{1}{2n}\boldsymbol{1}_{2n}^{\top}\left(\boldsymbol{y}_{T}-\boldsymbol{y}_{C}\right)=\tau~{}~{}\blacksquare$
@@
$\mathbb{V}\text{ar}\left[W_{i}\right]$ $=$ $1-b=\frac{n_{T}n_{C}}{4n^{2}}=r\tilde{r}$ (25)
@@ A.7 Variance of the MSE
$=\sum_{i=1}^{2n}\sigma_{i\,,i}^{2}\mathbb{E}\left[Z_{i}^{4}\right]+\sum_{i=j\neq k=\ell}\sigma_{i\,,i}\sigma_{k\,,k}\rho_{i}\rho_{k}+\sum_{i=k\neq j=\ell}\sigma_{i\,,j}\sigma_{i\,,j}\rho_{i}\rho_{j}+\sum_{i=\ell\neq j=k}\sigma_{i\,,j}\sigma_{j\,,i}\rho_{i}\rho_{j}.$
@@
$=$ $\boldsymbol{r}^{2}\tilde{r}^{2}\sum_{i=1}^{2n}\mathbb{E}\left[Z_{i}^{4}\right]+r^{2}\tilde{r}^{2}\sum_{i=j\neq k=\ell}\rho_{i}\rho_{k}+\sum_{i=k\neq j=\ell}\sigma_{i\,,j}^{2}\rho_{i}\rho_{j}+\sum_{i=\ell\neq j=k}\sigma_{i\,,j}^{2}\rho_{i}\rho_{j}$ (35)
$=$ $r^{2}\tilde{r}^{2}\left(\sum_{i=1}^{2n}\mathbb{E}\left[Z_{i}^{4}\right]+\sum_{i\neq j}\rho_{i}\rho_{j}\right)+2\sum_{i\neq j}\sigma_{i\,,j}^{2}\rho_{i}\rho_{j}.$
@@
$[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}]_{i,i}=\left[\sigma_{i\,,1}\rho_{1}\ldots\sigma_{i\,,2n}\rho_{2n}\right]\left[\sigma_{1\,,j}\rho_{j}\ldots\sigma_{2n\,,j}\rho_{j}\right]^{\top}=\sum_{j=1}^{2n}(\sigma_{i\,,j}\rho_{j})(\sigma_{j\,,i}\rho_{i})=\sum_{j=1}^{2n}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}$
@@
$\footnotesize 2\,\text{tr}\left[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}\right]=2\sum_{i=1}^{2n}\sum_{j=1}^{2n}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}=2\sum_{i\neq j}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}+2\sum_{i=j}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}=2\sum_{i\neq j}\sigma_{i\,,j}^{2}\rho_{j}\rho_{i}+2r^{2}\tilde{r}^{2}\sum_{i=1}^{2n}\rho_{i}^{2}$ (36)
@@
$\mathbb{E}\left[(\boldsymbol{Z}^{\top}\boldsymbol{\Sigma}_{W}\boldsymbol{Z})^{2}\right]=r^{2}\tilde{r}^{2}\left(\sum_{i=1}^{2n}\mathbb{E}\left[Z_{i}^{4}\right]+\sum_{i\neq j}\rho_{i}\rho_{j}-2\sum_{i=1}^{2n}\rho_{i}^{2}\right)+2\,\text{tr}\left[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}\right].$
@@
$\mathbb{V}\text{ar}_{\boldsymbol{Z}}\left[\boldsymbol{Z}^{\top}\boldsymbol{\Sigma}_{W}\boldsymbol{Z}\right]=r^{2}\tilde{r}^{2}\left(\sum_{i=1}^{2n}\mathbb{E}\left[Z_{i}^{4}\right]+\sum_{i\neq j}\rho_{i}\rho_{j}-2\sum_{i=1}^{2n}\rho_{i}^{2}-\text{tr}\left[\boldsymbol{\Sigma}_{Z}\right]^{2}\right)+2\,\text{tr}\left[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}\right].$
Substituting the expansion of $\text{tr}\left[\boldsymbol{\Sigma}_{Z}\right]^{2}=\sum_{i\neq j}\rho_{i}\rho_{j}+\sum_{i=1}^{2n}\rho_{i}^{2}$ we arrive at
$\mathbb{V}\text{ar}_{\boldsymbol{Z}}\left[\boldsymbol{Z}^{\top}\boldsymbol{\Sigma}_{W}\boldsymbol{Z}\right]$ $=$ $r^{2}\tilde{r}^{2}\left(\sum_{i=1}^{2n}\mathbb{E}\left[Z_{i}^{4}\right]-3\sum_{i=1}^{2n}\rho_{i}^{2}\right)+2\,\text{tr}\left[(\boldsymbol{\Sigma}_{W}\boldsymbol{\Sigma}_{Z})^{2}\right]$ (37)